OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y =-2x^2 - x - 12. A. -48 B. -12 C. -95/8 D. 11

OpenStudy (anonymous):

Take derivative, -4 x -1 and set = 0 x = -1/4 evaluate f(-1/4) = -95/8 Second derivative, -4 ,says slope is decreasing, thus this is a maximum.

OpenStudy (anonymous):

thanks